Find the soultion to the system of equations.
x = -y
y + x = 6

Answer by Alan3354(21581)

Find the soultion to the system of equations.
x = -y --> x + y = 0
y + x = 6
y + x = 0
-----------
The equations are inconsistent, no solution.